OpenStudy (babygal0831):

how do you solve for Y-1=-5(X-2)

OpenStudy (welshfella):

when you say solve i am assuming you want to find Y in terms of x? Y = something?

OpenStudy (welshfella):

I'm talking to myself again!!

OpenStudy (giitter):

Simplifying y + -1 = -5(x + -2) Reorder the terms: -1 + y = -5(x + -2) Reorder the terms: -1 + y = -5(-2 + x) -1 + y = (-2 * -5 + x * -5) -1 + y = (10 + -5x) Solving -1 + y = 10 + -5x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + y = 10 + 1 + -5x Combine like terms: -1 + 1 = 0 0 + y = 10 + 1 + -5x y = 10 + 1 + -5x Combine like terms: 10 + 1 = 11 y = 11 + -5x Simplifying y = 11 + -5x

OpenStudy (wanell):

(y - 1) - (0 - 5 • (x - 2)) = 0 Step 2 : Equation at the end of step 2 : y + 5x - 11 = 0 Step 3 : Equation of a Straight Line 3.1 Solve y+5x-11 = 0

OpenStudy (wanell):

1. Slope = -10.000/2.000 = -5.000 2. x-intercept = 11/5 = 2.20000 3. y-intercept = 11/1 = 11.00000